p=<3,5>;q=<2,4>
Find the angle between the vectors to the nearest tenth of a degree.
Select one:
a. 4.4°
b. 64.6°
c. 5.4°
d. 63.4°

Answer by Alan3354(69443)

p=<3,5>;q=<2,4>
Find the angle between the vectors to the nearest tenth of a degree.
------------
p.q = |p|*|q|*cos(A) --- dot product
p.q = 3*2 + 4*5 = 26
---
|p| = sqrt(3^2 + 5^2) = sqrt(34)
|q| = sqrt(2^2 + 4^2) = 2sqrt(5)
-----------
26 = 2sqrt(170)*cos(A)
cos(A) =~ 0.99705
A =~ 4.4 degs